OpenStudy (acacia):

Find f o g if f(x) = x + 7 and g(x) = 3x – 2

OpenStudy (anonymous):

Gah typed an incorrect answer; f o g = f(g(x) in other words substitute g(x) the function when you see x in f(x) the function

OpenStudy (anonymous):

To fine f of g just place g (3x -2 ) into f(x) so f(x) = 3x -2 + 7 now solve for f(x)

OpenStudy (anonymous):

\[f(g(x)) = (3x-2)+7 = 3x+5\]

OpenStudy (acacia):

3x+5